OpenStudy (codym251):

drag the fractions to put them in order from least to greatest. 2/3 1/5 5/6 9/10

OpenStudy (aveline):

1/5 is the smallest and 9/10 is the largest. Can you figure out the sequence?

OpenStudy (codym251):

what about 2/3 and 5/6

OpenStudy (aveline):

Right

OpenStudy (codym251):

which one comes next after 1/5?

OpenStudy (aveline):

Which is larger, 2/3 or 5/6? Try changing the denominator of 2/3 if you can't figure it out.

OpenStudy (codym251):

i thinck 2/3 is the largest

OpenStudy (codym251):

i mean think

OpenStudy (aveline):

\[\frac{ 2 }{ 3 }=\frac{ ? }{ 6 }\]

OpenStudy (aveline):

@codym251

OpenStudy (codym251):

yes

OpenStudy (aveline):

Solve the equation

OpenStudy (codym251):

i am trying to

OpenStudy (codym251):

2/3 is the largest

OpenStudy (aveline):

\[\frac{ 2 }{ 3 } \times \frac{ 2 }{ 2 }=?\]

OpenStudy (codym251):

4/8 right?

OpenStudy (aveline):

Try again

OpenStudy (codym251):

how do you do it then?

OpenStudy (aveline):

You multiplied incorrectly.

OpenStudy (codym251):

do crisscross then

OpenStudy (aveline):

._." What is 2 times 3?

OpenStudy (codym251):

6

OpenStudy (aveline):

\[\frac{ 2 }{ 3 } \times \frac{ 2 }{ 2 }=?\]

OpenStudy (codym251):

4/6 right

OpenStudy (aveline):

Correct. Now can you tell which, 4/6 or 5/6, is larger?

OpenStudy (codym251):

yes thank you

OpenStudy (codym251):

i have a 7 more to do. can you help me?

OpenStudy (codym251):

see you tomorrow Aveline

OpenStudy (aveline):

If you need more help, open a new question and tag me.

OpenStudy (codym251):

ok

OpenStudy (codym251):

can i close this question now?

OpenStudy (codym251):

can i

OpenStudy (aveline):

yes

OpenStudy (codym251):

ok

OpenStudy (codym251):

thanks for the help aveline